OpenStudy (anonymous):

The coefficient of friction opposing the sliding of a 5.00 kg rubber chicken across the top of a table is 0.408. How hard must you pull to make the rubber chicken slide with an acceleration of 2.00 m/s2?

OpenStudy (anonymous):

I assume that coefficient of friction is for kinetic friction, and no static friction? Anyway, use the formula\[F=ma\]m is the mass of the chicken, and a will be 2.00 F is going to be how hard you're pulling minus the friction force:\[F=Q-N \times \mu_{k}\]So you want to solve\[P-N \times \mu_{k}=m \times a\](N is the normal force, which is equal and opposite of gravity if the object is not moving vertically on the table).

OpenStudy (anonymous):

That Q should be a P (though it honesty doesn't matter what you call it).

OpenStudy (anonymous):

I have this, (roughly equivalent to yours)... is this correct?\[F _{NET}=ma\] \[(F _{NET}=F _{Friction})=ma\] \[(\mu F _{NET})=MA\]

OpenStudy (anonymous):

(.408mu*9.8)=ma

OpenStudy (anonymous):

What exactly are we solving for?

OpenStudy (anonymous):

Not quite. You have\[F_{Friction}=ma\]But I have\[F_{You}-F_{Friction}=ma\]and you are solving for \[F_{You}\]

OpenStudy (anonymous):

By (μFNET)=MA I meant (μFN)=ma

OpenStudy (anonymous):

oh,okay!!

OpenStudy (anonymous):

So could we say that Fyou= tension?

OpenStudy (anonymous):

I suppose if you had a rope or string attached to the rubber chicken, though it isn't needed. Why do you ask?

OpenStudy (anonymous):

Oh, in the diagram it has a string attached, and I just had to label the force

OpenStudy (anonymous):

Ahh, ok. The only thing is that this was more of a physics question, so in future you might want to ask similar questions in that forum instead.

OpenStudy (anonymous):

Noted! Thanks a million by the way!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

One more thing, would the value we're trying to find be acceleration or velocity?

OpenStudy (anonymous):

Oh wait it would be force, wouldn't it

OpenStudy (anonymous):

Yep! :P

OpenStudy (anonymous):

:)